OpenStudy (anonymous):

for these three equations x^2-4x+3=0 x^2-7x+2=20 x^2-3x=1 solve at least one by factoring solve at least one by using the quadratic formula solve the last one with either of the two options

OpenStudy (anonymous):

solve the second one by factoring and the last one by factoring also the first one you need to solve with quad eq.

OpenStudy (anonymous):

First of all they all need to = 0. The first two should be factoring.. the last one is quadratic formula.

OpenStudy (anonymous):

x^2 - 4x + 3 = 0 Hint (3 * 1 = 3) (3 + 1 = 4)

OpenStudy (anonymous):

1. (x - 3)(x - 1)

OpenStudy (anonymous):

yea hes right i copied in diff order

OpenStudy (anonymous):

x^2 - 7x + 2 = 20 x^2 - 7x - 18 = 0 Factors of 18 that subtract to give you 7 What are they?

OpenStudy (anonymous):

x^2 - 3x - 1 = 0 the reason you use the quadratic formula is because there are no factors of 1 that subtract to give you 3.

OpenStudy (anonymous):

Confused.. what are the solutions to #1 if (x - 3)(x - 1) = 0

OpenStudy (anonymous):

3 and 1????

OpenStudy (anonymous):

Very good

OpenStudy (anonymous):

Now on the 2nd one... the factors of 18 are 1 * 18 2 * 9 3 * 6 which pair subtracts to give you 7

OpenStudy (anonymous):

2 and 9

OpenStudy (anonymous):

good so you have (x - 9)(x + 2) = 0 what would be your answers for x

OpenStudy (anonymous):

9 and-2

OpenStudy (anonymous):

Very good. Now we need to do the quadratic formula.. do you know what that is? yes/no

OpenStudy (anonymous):

i think so ax^2+bx+c

OpenStudy (anonymous):

well.. you have to get it in this form first.. so we have x^2 - 3x - 1 = 0 what are a? b? c?

OpenStudy (anonymous):

a=1 b=-3 c=-1

OpenStudy (anonymous):

Good. the formula is x = -b +- sqrt(b^2 - 4ac) all over 2a Have you seen this the +- means a + with a - sign under it.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

-b I like to read as the opposite of b so the opposite of -3 is +3

OpenStudy (anonymous):

x = 3 +- sqrt((-3)^2 - 4(1)(-1) ------------------------ 2(1)

OpenStudy (anonymous):

x = 3 +- sqrt(9 +4) ------------- 2 do you see how I got this? the (-4)(1)(-1) is a + 4

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

OK.. so your answer is x = 3 +- sqrt(13) ----------- 2

OpenStudy (anonymous):

so 4.80 and 1.20

OpenStudy (anonymous):

I don't think so..

OpenStudy (anonymous):

If you are using a calculator put in (3 + sqrt(13))/2 (3 - sqrt(13))/2

OpenStudy (anonymous):

3.307

OpenStudy (anonymous):

-3.027

OpenStudy (anonymous):

Oops.. -.3027

myininaya (myininaya):

but those aren't the answers those are approximations since the sqrt(13) is irrational